Express the complex number in trigonometric form.

-2

Help me please
1
Expert's answer
2014-05-14T13:31:19-0400
-2=|z|(cos(alpha)+i*sin(alpha))=2(cos(pi)+isin(pi)) is the trigonometric form of this complex number (the answer is the same for cases when a polar angle ranges from -pi to pi and when a polar angle ranges from 0 to 2*pi).